Two doctors are known to agree 90% when doing independent assessments of patients. If they assess 6 patients, the probability they will agree exactly 4 times is:

Select one:

a. 0.0639

b. 0.0287

c. 0.221

d. 0.098

e. 0.0762
Answer #1

Here, n = 6, p = 0.9, (1 - p) = 0.1 and x = 4
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 4)
P(X = 4) = 6C4 * 0.9^4 * 0.1^2
P(X = 4) = 0.098
0